OpenStudy (anonymous):

How would I find the x and y intercept for f(x)=-1/(0-2)^2? I know I have to make x=0 for y and y=0 for x, but I'm not sure I know how to solve it.

OpenStudy (anonymous):

That zero that's already there is an x right?

OpenStudy (anonymous):

Oh yeah my bad, yes it's supposed to be a x. So the equation is f(x)=-1/(x-2)^2. Sorry lol.

OpenStudy (anonymous):

Lol no you're fine! So yep, for y-intercept, you plug in 0 for x. So y=-1/(0-2)^2 Order of operations say you do what's in parentheses first, then exponents. So, y=-1/(-2)^2 So y= -1/4 Are you good so far?

OpenStudy (anonymous):

Yeah I am. And then next would I bring 4 to the other side by multiplying?

OpenStudy (anonymous):

Nope, you've already got y by itself, so -1/4, or (0, -1/4) is the y-intercept :)

OpenStudy (anonymous):

Ohhh, yeah that makes sense. Lol I'm so out of it today, I'm wayyyyy to tired.

OpenStudy (anonymous):

But to find x is confusing me. Like I know I'd set it up as 0=-1/(x-2)^2. But from there I'm confused.

OpenStudy (anonymous):

Oh jeez, that is weird So (x-2)^2 is the same as (x-2)(x-2) So you'd FOIL/clam shell/whatever your method is for that first

OpenStudy (anonymous):

Should I say it, and use it as an example or do you wanna try first lol

OpenStudy (anonymous):

Let me try it first real quick, I needa practice anyways :)

OpenStudy (anonymous):

I figured! I'm an answer-blurter, so I stopped myself hahaha

OpenStudy (anonymous):

Ok so I got 0=-1/x^2-4x+2

OpenStudy (anonymous):

Eep, so close, the last number is 4, since -2 times -2 is 4. Shizzzz you didn't even need to do that though now that I look at the overall answer. Good practice though, dont kill me ;) But what they wanted you to do was flip the fraction (with (x-2)^2 from the bottom to the top) That makes the whole thing positive so it'll look like 0=(x-2)^2 /1 Then square root everything like \[\sqrt{0}=\sqrt{(x-2)^{2}}\] Which leaves 0= + or - (x-2) Moving the two over makes x= 2 and -2 ^So those are you intercepts. Ugh sorry for the confusion at first, I hope this helps!!

OpenStudy (anonymous):

Oh right! Lol I must be really tired, I did get 4 at first I just wrote down 2. Ugh, school just started so my brain isn't used to this yet. But ok, that makes sense. I just don't know why you'd flip it.

OpenStudy (anonymous):

Well if you left it how it was, like 0=-1/(x-2)^2, the only way to try and solve it would be multiplying both sides by the denominator to cancel it out. But that would leave 0=-1, which is not true. So when things like that happen, the answer would be undefined. BUT in this case, we could flip it, and still come out with x intercepts. Usually undefined as an answer for intercepts doesnt make sense anyway. Does that make sense or am I just frying your brain more? Lol

OpenStudy (anonymous):

Ok yeah that makes sense. I just didn't know you were allowed to flip it and all.

OpenStudy (anonymous):

Yep! You can flip fractions, but when you do, the whole top and bottom need to switch. And if it was a positive fraction, it becomes negative. Negative fractions become positive.

OpenStudy (anonymous):

It's called reciprocals if you wanna google for later or something

OpenStudy (anonymous):

Oh yeah I remember, ok yeah that makes sense. Thanks so much :) btw quick question, are you good at rational functions? With the whole asymptotes and all that? I'm in precal btw.

OpenStudy (anonymous):

Ugh pre-cal was my worst math year tbh. I was failing until like the very end when we got extra credit and exempted assignments for random crap lol so I probably wouldn't be the ebst person to ask :/

OpenStudy (anonymous):

Oh that's fine, thanks for you help tho. I really appreciate it, I wouldn't have been able to do my hw otherwise.

OpenStudy (anonymous):

Haha aw no problem! Good luck!